# The Surprising Effectiveness of Negative Reinforcement in LLM Reasoning

### v1

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) is a promising approach for training language models (LMs) on reasoning tasks that elicit emergent long chains of thought (CoTs). Unlike supervised learning, it updates the model using both correct and incorrect samples via policy gradients. To better understand its mechanism, we decompose the learning signal into reinforcing correct responses and penalizing incorrect ones, referred to as Positive and Negative Sample Reinforcement (PSR and NSR), respectively. We train Qwen2.5-Math-7B and Qwen3-4B on a mathematical reasoning dataset and uncover a surprising result: training with only negative samples—without reinforcing correct responses—can be highly effective: it consistently improves performance over the base model across the entire Pass@$k$ spectrum ($k$ up to $256$), often matching or surpassing PPO and GRPO. In contrast, reinforcing only correct responses improves Pass@$1$ but degrades performance at higher $k$, due to reduced diversity. These inference-scaling trends highlight that solely penalizing incorrect responses may contribute more to performance than previously recognized. Through gradient analysis, we show that NSR works by suppressing incorrect generations and redistributing probability mass toward other plausible candidates, guided by the model’s prior beliefs. It refines the model’s existing knowledge rather than introducing entirely new behaviors. Building on this insight, we propose a simple variant of the RL objective that upweights NSR, and show that it consistently improves overall Pass@$k$ performance on MATH, AIME 2025, and AMC23. Our code is available at https://github.com/TianHongZXY/RLVR-Decomposed.
@@ 1 Introduction
Language models (LMs) have recently demonstrated remarkable capabilities in various complex reasoning tasks, including mathematics cobbe2021training ; hendrycks2021measuring , coding jimenezswe ; yang2024swe , and scientific reasoning phan2025humanity ; rein2024gpqa . A key technique in achieving such success is reinforcement learning with verifiable rewards (RLVR) guo2025deepseek ; jaech2024openai ; lambert2024t ; team2025kimi , which is particularly effective in domains where the correctness of an outcome can be automatically verified via tools or functions. RLVR typically employs a binary reward ($+1$ or $-1$) based on the objective correctness of model responses. This simple yet effective mechanism not only mitigates reward hacking miaoinform ; skalse2022defining but also eliminates the need for extensive human annotations and complex reward model training li2025process ; Zhang2025TheLO .
RLVR’s appeal is multifaceted: it offers a conceptually simple formulation lambert2024t , exhibits notable sample efficiency fatemi2025concise ; li2025limr ; wang2025reinforcement , and enables inference-time scaling behaviors gandhi2025cognitive ; muennighoff2025s1 ; wu2025more ; yeo2025demystifying ; zeng2025simplerl . However, the precise mechanisms driving its effectiveness remain underexplored, particularly how it utilizes correct and incorrect samples. To address this, we decompose RLVR into two learning paradigms: Positive Sample Reinforcement (PSR) and Negative Sample Reinforcement (NSR), as illustrated in Figure 1. This decomposition prompts a natural question: What roles do PSR and NSR play in shaping model behavior and generalization?
@@ 2.1 Reinforcement Learning with Verifiable Rewards
Reinforcement learning with verifiable rewards (RLVR) has recently emerged as a powerful paradigm for enabling LMs to self-improve on tasks with objectively verifiable outcomes. The reward is given by a deterministic verification function $r$ which assesses whether the model’s response $\boldsymbol{y}$ to the prompt $\boldsymbol{x}$ is correct or not. All tokens $\{y_{1},y_{2},\dots,y_{T}\}$ in a response $\boldsymbol{y}$ receive the same reward (e.g., $+1$ for correct responses and $-1$ for incorrect ones).
@@ 3.1 Experimental Setup
To understand how different training objectives affect the model’s behavior, we train the model with PSR and NSR and evaluate their inference scaling performance on reasoning tasks. Specifically, we use Qwen2.5-Math-7B (qwen2.5_report, ) and Qwen3-4B (yang2025qwen3, ) as the base models. Qwen3-4B has two modes (thinking and non-thinking), and we use the non-thinking mode for training and inference.
@@
We compare the performance of PSR and NSR with commonly used RL algorithms, including PPO schulman2017proximal and GRPO guo2025deepseek ; shao2024deepseekmath . PSR and NSR are implemented by selectively updating the policy model using only correct or incorrect responses, respectively. As a result, PSR and NSR are trained on fewer samples per batch than standard RL algorithms (e.g., PPO and GRPO) that use both correct and incorrect responses. The training objectives of these algorithms can be found in Section C.1. We also report the performance of the base models for reference.
@@
For the training set, we use MATH (hendrycks2021measuring, ), which contains $7{,}500$ problems. We train the models using the verl framework (sheng2024hybridflow, ). The prompt batch size is $1{,}024$, with $8$ rollouts generated per prompt. The sampling temperature during training is set to $1.0$, and the maximum context length is set to $4{,}096$ and $32{,}768$ tokens for Qwen2.5-Math-7B and Qwen3-4B, respectively. We update the model with a mini-batch size of $256$ and a learning rate of 1e-6. More hyperparameter settings can be found in Section C.1.
@@
We evaluate on three widely used math reasoning benchmarks, including the test sets of MATH, AIME 2025 and AMC23. During evaluation, we sample $256$ responses per prompt for Qwen2.5-Math-7B with a temperature of $0.6$ and a top-$p$ of $0.95$, and $64$ responses for Qwen3-4B with a temperature of $0.7$, a top-$p$ of $0.8$ and a top-$k$ of $20$, prompt templates can be found in Appendix C.2. For evaluation metric, recent work hochlehnert2025sober highlights that accuracy based on greedy decoding can be unreliable. For this reason, we adopt a full spectrum of Pass@$k$ as our main evaluation metric, using $k\in\{1,2,4,8,16,32,64,128,256\}$ for Qwen2.5-Math-7B and $k\in\{1,2,4,8,16,32,64\}$ for Qwen3-4B. Pass@$k$ is defined as the fraction of problems for which at least one correct response is produced in $k$ independent trials. However, directly computing Pass@$k$ using only $k$ samples per example often suffers from high variance. We follow the unbiased estimator proposed by chen2021evaluate_llm_on_code , which generates $n$ samples per problem ($n\geq k$), counts the number of correct responses $c$, and computes an unbiased estimate of Pass@$k$ as:
[…]
@@ 3.2 Inference Scaling Trends Under Different Training Objectives
As shown in Figures 2 and 3, NSR exhibits unexpectedly strong performance across the full range of $k$ values. Despite being trained solely on negative samples, it consistently improves Pass@$k$ compared to the base model. While PPO, GRPO, and PSR explicitly reinforce correct responses and naturally outperform the base model at Pass@$1$, it is surprising that NSR achieves a comparable Pass@$1$ without training on any correct responses. This suggests that NSR is able to reinforce correct responses indirectly by suppressing incorrect ones and redistributing probability mass toward plausible alternatives.
@@
At larger decoding budgets (e.g., Pass@$256$), recent work (yue2025does, ) shows that RL-trained models often lose their advantage, and in some cases underperform the base model. This trend is generally observed in our experiments with PPO, GRPO, and PSR, especially in Figure 2. NSR, on the other hand, maintains a comparable or even better Pass@$256$ performance than the base model, generally outperforming the other algorithms. These results suggest that NSR promotes exploration and preserves the output diversity.
@@
In contrast, PSR, which only reinforces correct samples, displays a more polarized behavior. It improves Pass@$1$, particularly on AIME 2025 and AMC23 in Figure 2. However, this precision comes at a cost: as $k$ increases, Pass@$k$ improves more slowly than other methods and eventually falls below the base model for $k>8$. As shown in Figure 4(a), PSR improves greedy decoding accuracy rapidly during early training but plateaus quickly and is overtaken by other methods. This behavior indicates that PSR overly concentrates probability mass on early correct responses, leading to overconfidence and a collapsed output distribution and ultimately limiting the model’s ability to generate diverse correct responses when allowing for more test-time compute.
@@
Figure 3 demonstrates the performance of training Qwen3-4B under non-thinking mode to simulate a scenario where the model’s underlying knowledge is known to be strong (i.e., its learned thinking mode triggered by the ‘<think>’ tag). While one may intuitively expect all algorithms to easily transfer the model’s reasoning ability across different prompt formats (i.e., from thinking to non-thinking mode), our results demonstrate significant performance differences across different algorithms. PSR fails to activate these latent capabilities and does not improve the performance, even degrading Pass@$k$ on MATH and AMC23 significantly. This highlights a fundamental limitation of PSR: it reinforces the currently dominant behavior in the output distribution, suppressing potentially stronger underlying capabilities. In contrast, NSR and GRPO significantly improve the performance of the base model across all Pass@$k$ metrics, achieving thinking mode capabilities. Specifically, Qwen3-4B in thinking mode achieves a Pass@$1$ of $94.5$ and a Pass@$64$ of $97.8$ on MATH test set. NSR and GRPO closely match this performance, with NSR reaching $94.0$ (Pass@$1$) and $98.0$ (Pass@$64$), and GRPO achieving $93.9$ (Pass@$1$) and $98.2$ (Pass@$64$).
@@ 4.1 Entropy as a Lens on Inference Scaling Performance
Since diversity is crucial for strong Pass@$k$ performance—especially at a large $k$—we quantify model diversity by tracking its entropy on a held-out test set throughout training, aiming to understand how entropy evolves under different training algorithms. In addition, we monitor two complementary metrics on the training set: the correct sample ratio, which measures the proportion of correct responses per batch, and the fully solved ratio, which measures the fraction of problems per batch where all rollouts are correct.
@@
As shown in Figures 4(c) and 4(d), NSR consistently improves the correct sample ratio but maintains a lower proportion of correct samples and fully solved problems throughout training. This indicates that the model avoids becoming overconfident in the observed correct responses. By preserving uncertainty, NSR enables stronger scaling performance in Pass@$k$, as demonstrated in Figure 2. In contrast, PSR rapidly increases the number of correct samples in each batch and achieves a higher fully solved ratio compared to other methods, suggesting a tendency to overfit to the observed correct responses. While this leads to better Pass@$1$ performance, it significantly reduces output diversity and results in weaker Pass@$k$ performance as $k$ increases.
@@ 4.2 Token-Level Gradient Dynamics of PSR and NSR
Figure 5: Gradient dynamics of PSR and NSR under a math word problem example. Bars indicate token logits before the update, and arrows indicate gradient direction and magnitude. Left (PSR): The model generates the correct response (“7”) and receives a $+1$ reward. Gradients increase the logit of the sampled token while suppressing all others, including potentially correct alternatives like “2”, resulting in a sharpened, overconfident distribution. Right (NSR): The model generates an incorrect response (“6”) and receives a $-1$ reward. Gradients demote the sampled incorrect token and proportionally reallocate logits to other tokens (e.g., “7”, “2”) based on their current probabilities, thereby promoting exploration on alternative correct paths and preserving diversity.
To analyze the effect of these objectives on the model’s token distribution, we compute the gradient of the loss with respect to token-level logits at each step. Let $\pi_{v}=\pi_{\theta}(v|\boldsymbol{x},\boldsymbol{y}_{<t})$ denote the probability of token $v$ at time step $t$, the gradient descent directions of PSR and NSR are shown in Equations 7 and 8 (the full derivation is provided in Appendix A), which are illustrated in Figure 5.
$-\frac{\partial\mathcal{L}_{\text{PSR}}}{\partial z_{v}}\propto\begin{cases}\pi_{y_{t}}\cdot(1-\pi_{y_{t}})&\text{if }v=y_{t}\quad\text{(sampled token)}\\ -\pi_{v}\cdot\pi_{y_{t}}&\text{if }v\neq y_{t}\quad\text{(unsampled token)}\end{cases}$ (7)
@@
These analyses highlight an important strength of NSR: it preserves the model’s prior knowledge over plausible tokens and stops updating once errors are corrected, effectively locking in successful experiences. This is a unique advantage of NSR, as we demonstrate in Appendix A via gradient analysis that simply applying an entropy bonus in the policy loss, despite promoting diversity, cannot preserve model’s prior.

### v2

@@ Abstract
Reinforcement learning with verifiable rewards (RLVR) is a promising approach for training language models (LMs) on reasoning tasks that elicit emergent long chains of thought (CoTs). Unlike supervised learning, it updates the model using both correct and incorrect samples via policy gradients. To better understand its mechanism, we decompose the learning signal into reinforcing correct responses and penalizing incorrect ones, referred to as Positive and Negative Sample Reinforcement (PSR and NSR), respectively. We train Qwen2.5-Math-7B, Qwen3-4B and Llama-3.1-8B-Instruct on a mathematical reasoning dataset and uncover a surprising result: training with only negative samples—without reinforcing correct responses—can be highly effective: it consistently improves performance over the base model across the entire Pass@$k$ spectrum ($k$ up to $256$), often matching or surpassing PPO and GRPO. In contrast, reinforcing only correct responses improves Pass@$1$ but degrades performance at higher $k$, due to reduced diversity. These inference-scaling trends highlight that solely penalizing incorrect responses may contribute more to performance than previously recognized. Through gradient analysis, we show that NSR works by suppressing incorrect generations and redistributing probability mass toward other plausible candidates, guided by the model’s prior beliefs. It refines the model’s existing knowledge rather than introducing entirely new behaviors. Building on this insight, we propose a simple variant of the RL objective that upweights NSR, and show that it consistently improves overall Pass@$k$ performance on MATH, AIME 2025, and AMC23. Our code is available at https://github.com/TianHongZXY/RLVR-Decomposed.
@@ 1 Introduction
Language models (LMs) have recently demonstrated remarkable capabilities in various complex reasoning tasks, including mathematics cobbe2021training ; hendrycks2021measuring , coding jimenezswe ; yang2024swe , and scientific reasoning phan2025humanity ; rein2024gpqa . A key technique in achieving such success is reinforcement learning with verifiable rewards (RLVR) guo2025deepseek ; jaech2024openai ; lambert2024t ; team2025kimi , which is particularly effective in domains where the correctness of an outcome can be automatically verified via tools or functions. RLVR typically employs a binary reward (+1+1 or 1-1) based on the objective correctness of model responses. This simple yet effective mechanism not only mitigates reward hacking miaoinform ; skalse2022defining but also eliminates the need for extensive human annotations and complex reward model training li2025process ; Zhang2025TheLO .
RLVR’s appeal is multifaceted: it offers a conceptually simple formulation lambert2024t , exhibits notable sample efficiency fatemi2025concise ; li2025limr ; wang2025reinforcement , and enables inference-time scaling behaviors gandhi2025cognitive ; muennighoff2025s1 ; wu2025more ; yeo2025demystifying ; zeng2025simplerl . However, the precise mechanisms driving its effectiveness remain underexplored, particularly how it utilizes correct and incorrect samples. To address this, we decompose RLVR into two learning paradigms: Positive Sample Reinforcement (PSR) and Negative Sample Reinforcement (NSR), as illustrated in Figure˜1. This decomposition prompts a natural question: What roles do PSR and NSR play in shaping model behavior and generalization?
@@ 2.1 Reinforcement Learning with Verifiable Rewards
Reinforcement learning with verifiable rewards (RLVR) has recently emerged as a powerful paradigm for enabling LMs to self-improve on tasks with objectively verifiable outcomes. The reward is given by a deterministic verification function $r$ which assesses whether the model’s response $\boldsymbol{y}$ to the prompt $\boldsymbol{x}$ is correct or not. All tokens $\{y_{1},y_{2},\dots,y_{T}\}$ in a response $\boldsymbol{y}$ receive the same reward (e.g., +1+1 for correct responses and 1-1 for incorrect ones).
@@ 3.1 Experimental Setup
To understand how different training objectives affect the model’s behavior, we train the model with PSR and NSR and evaluate their inference scaling performance on reasoning tasks. Specifically, we use Qwen2.5-Math-7B (qwen2.5_report, ), Qwen3-4B (yang2025qwen3, ) and Llama-3.1-8B-Instruct (llama3, ) as the base models. Qwen3-4B has two modes (thinking and non-thinking), and we use the non-thinking mode for training and inference.
@@
We compare the performance of PSR and NSR with commonly used RL algorithms, including PPO schulman2017proximal and GRPO guo2025deepseek ; shao2024deepseekmath . PSR and NSR are implemented by selectively updating the policy model using only correct or incorrect responses, respectively. As a result, PSR and NSR are trained on fewer samples per batch than standard RL algorithms (e.g., PPO and GRPO) that use both correct and incorrect responses. The training objectives of these algorithms can be found in Section˜D.1. We also report the performance of the base models for reference.
@@
For the training set, we use MATH (hendrycks2021measuring, ), which contains 7,5007{,}500 problems. We train the models using the verl framework (sheng2024hybridflow, ). The prompt batch size is 1,0241{,}024, with $8$ rollouts generated per prompt. The sampling temperature during training is set to $1.0$, and the maximum context length is set to 4,0964{,}096, 4,0964{,}096 and 32,76832{,}768 tokens for Qwen2.5-Math-7B, Llama-3.1-8B-Instruct and Qwen3-4B, respectively. We update the model with a mini-batch size of $256$ and a learning rate of 1e-6. More hyperparameter settings can be found in Section˜D.1.
@@
We evaluate on three widely used math reasoning benchmarks, including the test sets of MATH, AIME 2025 and AMC23. During evaluation, we sample $256$ responses per prompt for Qwen2.5-Math-7B and Llama-3.1-8B-Instruct with a temperature of $0.6$ and a top-$p$ of $0.95$, and $64$ responses for Qwen3-4B with a temperature of $0.7$, a top-$p$ of $0.8$ and a top-$k$ of $20$, prompt templates can be found in Appendix D.2. For evaluation metric, recent work hochlehnert2025sober highlights that accuracy based on greedy decoding can be unreliable. For this reason, we adopt a full spectrum of Pass@$k$ as our main evaluation metric, using $k\in\{1,2,4,8,16,32,64,128,256\}$ for Qwen2.5-Math-7B and $k\in\{1,2,4,8,16,32,64\}$ for Qwen3-4B. Pass@$k$ is defined as the fraction of problems for which at least one correct response is produced in $k$ independent trials. However, directly computing Pass@$k$ using only $k$ samples per example often suffers from high variance. We follow the unbiased estimator proposed by chen2021evaluate_llm_on_code , which generates $n$ samples per problem ($n\geq k$), counts the number of correct responses $c$, and computes an unbiased estimate of Pass@$k$ as:
$\text{Pass@}k=\mathbb{E}_{\boldsymbol{x}\sim\mathcal{D}}\left[1-\frac{\binom{n-c}{k}}{\binom{n}{k}}\right$]. (5)
@@
Figure 4: Pass@$k$ curves of Llama-3.1-8B-Instruct trained with PPO, GRPO, PSR, and NSR. All the methods underperforms the base model, while NSR retains the most performances.
@@ 3.2 Inference Scaling Trends Under Different Training Objectives
As shown in Figures˜2 and 3, NSR exhibits unexpectedly strong performance across the full range of $k$ values. Despite being trained solely on negative samples, it consistently improves Pass@$k$ compared to the base model. While PPO, GRPO, and PSR explicitly reinforce correct responses and naturally outperform the base model at Pass@$1$, it is surprising that NSR achieves a comparable Pass@$1$ without training on any correct responses. This suggests that NSR is able to reinforce correct responses indirectly by suppressing incorrect ones and redistributing probability mass toward plausible alternatives.
@@
At larger decoding budgets (e.g., Pass@$256$), recent work (yue2025does, ) shows that RL-trained models often lose their advantage, and in some cases underperform the base model. This trend is generally observed in our experiments with PPO, GRPO, and PSR, especially in Figure˜2. NSR, on the other hand, maintains a comparable or even better Pass@$256$ performance than the base model, generally outperforming the other algorithms. These results suggest that NSR promotes exploration and preserves the output diversity.
@@
In contrast, PSR, which only reinforces correct samples, displays a more polarized behavior. It improves Pass@$1$, particularly on AIME 2025 and AMC23 in Figure˜2. However, this precision comes at a cost: as $k$ increases, Pass@$k$ improves more slowly than other methods and eventually falls below the base model for $k>8$. As shown in Figure 5(a), PSR improves greedy decoding accuracy rapidly during early training but plateaus quickly and is overtaken by other methods. This behavior indicates that PSR overly concentrates probability mass on early correct responses, leading to overconfidence and a collapsed output distribution and ultimately limiting the model’s ability to generate diverse correct responses when allowing for more test-time compute.
@@
Figure˜3 demonstrates the performance of training Qwen3-4B under non-thinking mode to simulate a scenario where the model’s underlying knowledge is known to be strong (i.e., its learned thinking mode triggered by the ‘<think>’ tag). While one may intuitively expect all algorithms to easily transfer the model’s reasoning ability across different prompt formats (i.e., from thinking to non-thinking mode), our results demonstrate significant performance differences across different algorithms. PSR fails to activate these latent capabilities and does not improve the performance, even degrading Pass@$k$ on MATH and AMC23 significantly. This highlights a fundamental limitation of PSR: it reinforces the currently dominant behavior in the output distribution, suppressing potentially stronger underlying capabilities. In contrast, NSR and GRPO significantly improve the performance of the base model across all Pass@$k$ metrics, achieving thinking mode capabilities. Specifically, Qwen3-4B in thinking mode achieves a Pass@$1$ of $94.5$ and a Pass@$64$ of $97.8$ on MATH test set. NSR and GRPO closely match this performance, with NSR reaching $94.0$ (Pass@11) and $98.0$ (Pass@$64$), and GRPO achieving $93.9$ (Pass@11) and $98.2$ (Pass@$64$).
RL training hurts Llama models’ inference scaling performance.
As shown in Figure˜4, applying different RL algorithms to Llama-3.1-8B-Instruct leads to degraded inference-scaling performance compared to the base model. Although Pass@$1$ improves on MATH and AMC23, Pass@$256$ drops substantially after RL training. This contrast between Qwen and Llama models highlights that backbone model itself often determines whether RL can bring benefits, consistent with recent works (yue2025does, ; octothinker, ; yeo2025demystifying, ). Among all methods, however, NSR causes the least degradation to the base model’s scaling behavior.
@@ 4.1 Entropy as a Lens on Inference Scaling Performance
Since diversity is crucial for strong Pass@$k$ performance—especially at a large kk—we quantify model diversity by tracking its entropy on a held-out test set throughout training, aiming to understand how entropy evolves under different training algorithms. In addition, we monitor two complementary metrics on the training set: the correct sample ratio, which measures the proportion of correct responses per batch, and the fully solved ratio, which measures the fraction of problems per batch where all rollouts are correct.
@@
As shown in Figures˜5(c) and 5(d), NSR consistently improves the correct sample ratio but maintains a lower proportion of correct samples and fully solved problems throughout training. This indicates that the model avoids becoming overconfident in the observed correct responses. By preserving uncertainty, NSR enables stronger scaling performance in Pass@$k$, as demonstrated in Figure 2. In contrast, PSR rapidly increases the number of correct samples in each batch and achieves a higher fully solved ratio compared to other methods, suggesting a tendency to overfit to the observed correct responses. While this leads to better Pass@$1$ performance, it significantly reduces output diversity and results in weaker Pass@$k$ performance as $k$ increases.
@@ 4.2 Token-Level Gradient Dynamics of PSR and NSR
Figure 6: Gradient dynamics of PSR and NSR under a math word problem example. Bars indicate token logits before the update, and arrows indicate gradient direction and magnitude. Left (PSR): The model generates the correct response (“7”) and receives a +1+1 reward. Gradients increase the logit of the sampled token while suppressing all others, including potentially correct alternatives like “2”, resulting in a sharpened, overconfident distribution. Right (NSR): The model generates an incorrect response (“6”) and receives a 1-1 reward. Gradients demote the sampled incorrect token and proportionally reallocate logits to other tokens (e.g., “7”, “2”) based on their current probabilities, thereby promoting exploration on alternative correct paths and preserving diversity.
To analyze the effect of these objectives on the model’s token distribution, we compute the gradient of the loss with respect to token-level logits at each step. Let $\pi_{v}=\pi_{\theta}(v|\boldsymbol{x},\boldsymbol{y}_{<t})$ denote the probability of token $v$ at time step $t$, the gradient descent directions of PSR and NSR are shown in Equations˜7 and 8 (the full derivation is provided in Appendix˜A), which are illustrated in Figure˜6.
[…]
@@
These analyses highlight an important strength of NSR: it preserves the model’s prior knowledge over plausible tokens and stops updating once errors are corrected, effectively locking in successful experiences. This is a unique advantage of NSR, as we demonstrate in Appendix B via gradient analysis that simply applying an entropy bonus in the policy loss, despite promoting diversity, cannot preserve model’s prior. We also include a comparison with the widely used unlikelihood training objective, a previous approach designed to suppress undesirable or incorrect generations (welleck2019neural, ; li2020don, ).